Sam has 2.05 in quarters and dimes. He has 4 more quarters than dimes. How many coins of each type does he have?
Answer by mananth(12270)

Sam has 2.05 in quarters and dimes. He has 4 more quarters than dimes. How many coins of each type does he have?
total value = 205 cents
..
let dimes be x
quarters = x+4
...
25(x+4)+10x=205
25x+100+10x=205
35x=205-100
35x=105
x= 3 dimes
quarters = 5
...
m.ananth@hotmail.ca